# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

### v1

@@ 3 Adversarial attacks against DNNs
The vulnerability of deep neural networks (DNNs) to adversarial samples has been first pointed out by Szegedy et al. in [1]. When the goal of the attacker is to induce a generic misclassification, we talk about an untargeted attack. When the misclassification is in favor of a specific class, we say that the attack is a targeted one. According to [1], the generation of a targeted adversarial example can be formally described as
$&\textrm{minimize}\quad||\delta||_{2}\\ &\textrm{s.t.}\quad\phi(x+\delta)=t\neq\phi(x)\\ &\textrm{and}\quad x+\delta\in[0,1]^{m}\end{aligned},$ (3)
@@
where $||\delta||_{2}$ denotes the $l_{2}$ norm of the perturbation $\delta$, and $t$ is the target class of the attack. The same formulation holds for the untargeted case, with the first constraint replaced by $\phi(x+\delta)\neq\phi(x)$. The constraint $x+\delta\in[0,1]^{m}$ makes sure that the resulting adversarial example is a valid input.
3.1 Basic adversarial attacks
Solving the minimization in (3) is generally hard, then the adversarial examples are determined by solving the simplified problem described below:
$\min_{\delta}$ $\hskip 5.69046pt\mathcal{L}({x+\delta},t)+\lambda\cdot\left\|\delta\right\|_{2}^{2}$ (4)
s.t. $\hskip 5.69046ptx+\delta\in[0,1]^{m}$
@@ 3.1 Basic adversarial attacks
where $\mathcal{L}({x+\delta},t)$ is the loss function under the target class $t$, and $\lambda$ is a parameter balancing the importance of the two terms. The problem in (4) can be solved by the box-constraint L-BFGS method [1]. Specifically, a line search is carried out to find the value of $\lambda>0$ for which the solution of (4) satisfies the adversarial condition, that is $\phi(x+\delta)=t$. Usually, the categorical cross entropy loss is considered. From (4), it is immediate to see that, in the one-hot case, when the categorical cross entropy loss is considered, only the target class contributes to the loss term, and then the attack only cares about increasing $f_{t}(x)$, regardless of the other scores, all the more that, due to the presence of the softmax, this also implies decreasing the other outputs.
To reduce the complexity of the L-BFGS attack, several suboptimal solutions have been proposed, that consider the problem - equivalent to (4) - of minimizing the loss function $\mathcal{L}$ subject to a constraint on the perturbation $\delta$. The Fast Sign Gradient Method (FGSM) method [2] obtains an adversarial perturbation in a computationally efficient way by considering the sign of the gradient of the output with respect to the input image. The FGSM attack is a one-step attack whose strength can be modulated only by increasing the value of the term multiplying the gradient sign. In [4], an iterative version of FGSM is introduced by applying FGSM multiple times, with a smaller step size, each time by recomputing the gradient. This method is often referred to as I-FGSM, or Basic Iterative Method (BIM). Another attack similar to the FGSM attack is the projected gradient descent (PGD) attack [5], that can be regarded to as a multi-step extension of the FGSM attack where the clip operation performed by BIM on the gradient (to force the solution to stay in the [0,1] range) is replaced by the gradient projection.
Other gradient-based attacks are the Jacobian-based Saliency Map Attack (JSMA) [3] and the DeepFool attack [6]. JSMA [3] consists of a greedy iterative procedure which relies on forward propagation to compute, at each iteration, a saliency map, indicating the pixels that contribute most to the classification. In such attack, the Jacobian matrix of the network function is used to compute the forward derivative and build the saliency map. A variant of this attack, called Maximal JSMA, has been proposed in [15]. DeepFool [6] is an efficient iterative attack that considers the minimal perturbation with respect to a linearized classifier. In this way, the current version of the adversarial example is projected onto the separating affine hyperplane at each iteration, until the classification boundary is crossed.
@@
The above gradient-based methods often fail due to the gradient vanishing problem [7]. A solution to overcome this issue has been proposed in [16], where the values of the logits are scaled by a temperature parameter. However, this approach lacks flexibility and the suitable distillation temperature parameter strongly depends on the network. A more flexible and efficient method working at the logits level has been proposed by Carlini and Wagner (C&W) [7]. In the C&W attack the loss term is replaced by a properly designed function depending on the difference between the logit of the target class and the largest among the logits of the other classes (since the logits do not sum one, two logits at a time must be considered). The perturbation is kept in the valid range by properly modifying the objective function. In this way, the following unconstrained problem is considered and solved (for the case of $L_{2}$ distortion metric):
$\min_{w}$ $\hskip 5.69046pt||\delta(w)||_{2}^{2}+$
@@
where $\delta(w)=1/2(\tanh(w)+1)-x$ is the distortion introduced in the image, and the constant parameter $\kappa>0$ is used to encourage the attack to generate a high confidence attacked image. Since logits are more sensitive to modifications of the input than the probability values obtained after the softmax activation C&W attack is less sensitive to vanishing gradient problems. However, we notice that, by considering only two logits at time, and neglecting the effect of the perturbation on the other logits, this approach is particularly suited to the one-hot case and is highly suboptimum when other output encoding schemes are used.
3.2 Attacks against DNNs with output encoding
@@
The adversarial attack algorithms described so far focus on models based on one-hot encoding. This is a favorable scenario for the attacker, who only needs to focus on increasing the output score (or logit) of the target class $t$. This is not the case with DNNs based on different output encoding schemes. The simplest examples are multi-label classifiers. From the attacker’s side, the difficulties of carrying out the attack in this case are due to the fact that the labels are not mutually exclusive, and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way. In addition, the attacker may want to modify more than one output label in a desired way. For instance, he may want to induce the classifier to interpret a daylight image showing a car driving in the rain, as a night image of a car driving with no rain. This prevents the application of the basic adversarial algorithms described in the previous section.
@@
The problem of crafting adversarial examples against multi-label classifiers has been addressed in [10], where the C&W and the DeepFool attacks have been extended to work in a multi-label setting. The extension of C&W attack, referred to as ML-C&W, works as follows: for every output node, a hinge loss term similar to the one in Eq. (3.1) is considered in the minimization. More specifically, the following term is added to the minimization: $\sum_{i=0}^{n}\max(0,\gamma-c_{ti}\cdot z_{i}(x+\delta))$, where $c_{t}$ is the target label vector and $\gamma$ is a confidence parameter. Note that the extension implicitly assumes that decoding is carried out element-wise. This is not true when all labels’ combinations are possible and with schemes adopting output channel coding. However, by working at the logits level, the C&W attack is general and can also be extended to work with different output encoding schemes (as discussed below).
@@
An extension of the DeepFool attack, referred to as ML-DF, has also been proposed in [10]. Instead of targeting an objective function minimizing the distortion, the method looks for the minimum perturbation that allows to enter the target region under a linear assumption on the model behavior. ML-DF works at the output score level, and hence is more prone to the vanishing gradient problem than ML-C&W. Given the threshold vector ${\rho}=({\rho}_{1},{\rho}_{2},..,{\rho}_{l})^{T}$ for the score level output ($\rho_{i}=0.5$ when $f_{i}\in[0,1]$, 00 when $f_{i}\in[-1,1]$), a vector with the distance to the boundary is derived and used to compute the perturbation. The perturbation obtained in this way is applied to the image. ML-DF applies the above procedure iteratively until the attack succeeds or a maximum number of iterations is reached. ML-DF is a greedy method and often leads to very large distortions; Moreover, the performance obtained are always inferior with respect to ML-C&W [10, 11].
@@
Inspired by ML-DF, another approach to implement an adversarial attack in the multi-label case, named Multi-Label Attack via Linear programming (MLA-LP), has been proposed in [11]. By assuming that the loss changes are linear for small distortions, a simplified formulation of the attack is solved to minimize the $L_{\infty}$ distortion introduced in the image subject to a constraint on the loss value, requiring that the final loss is lower than the threshold value. Thanks to the adoption of the $L_{\infty}$ distortion, the problem can be easily transformed into a linear programming problem, for which several solvers exist. Similarly to ML-DF, MLA-LP works at the final score output level. By considering the loss function instead of the model output function, MLA-LP tends to modify less the confidences of non-attacked labels with respect to ML-DF, with a lower distortion for the resulting attacked image.
@@
$\min_{\delta}$ $\hskip 5.69046pt\left(||\delta||_{2}^{2}-\lambda\cdot\min_{i}(c_{ti}\cdot z_{i}(x+\delta),\eta)\right),$ (6)
@@
3.2.1 The LOTS attack
A completely different kind of attack, which can be applied to any network regardless of the output encoding scheme is the Layerwise Origin-Target Synthesis (LOTS) attack [18]. LOTS generates the adversarial example by modifying the input sample so that its representation in the feature space is as similar as possible to that of a given target sample. Formally, given the internal representation $f^{-j}(x)$ of an image $x$ at layer $L-j$ and a target internal representation $f^{-j}(x_{t})$, LOTS attempts to minimize the term:
$||f^{-j}(x^{t})-f^{-j}(x)||^{2}.$ (7)
@@ 3.2.1 The LOTS attack
where $\nu^{-j}(x,x^{t})=\nabla_{x}\left(||f^{-j}(x^{t})-f^{-j}(x)||^{2}\right)$, and ||$\cdot$|| ⋅ | is applied element-wise, until the Euclidean distance between $f^{-l}(x)$ and the target is smaller than a predefined threshold. The final perturbed image $x^{adv}$ obtained in this way has an internal representation at layer $L-j$ that mimics that of the target sample $x^{t}$. In the LOTS algorithm, the target point $x^{t}$ is any point belonging to the target class of the attack. We observe that in general such $x^{t}$ may not be available, e.g. in the multilabel case when the attacker targets an arbitrary labels’ combination.
@@
3.2.2 JMA and prior art
Below we discuss the relationship and advantages of JMA with respect to the state-of-the-art attacks described above.
In contrast to LOTS [18], JMA defines the target point in the feature space that minimizes perturbation necessary to move the feature representation of the input sample to the target point. It does so, by relying on the Mahalanobis distance induced by the Jacobian matrix of the neural network function so to take into account the different effort, in terms of input perturbation, required to move the sample in different directions.
@@ 3.2.2 JMA and prior art
The ML-DF and MLA-LP methods [10, 11] are also based on the Jacobian matrix. However, in [10], the Jacobian matrix is used to implement a greedy algorithm, while in [11] is used to solve a suboptimum formulation of the attack. In our case, instead, we are able to solve the original formulation of the attack by Szegedy et al. under a first-order approximation. With such an approximation, in fact, the solution of the attack problem can be reduced to the solution of a constrained quadratic programming problem (where the objective function is a Mahalanobis distance term induced by the Jacobian). Moreover, our algorithm is a general one and can work either at the output level or at the logits level, regardless of the output encoding scheme used by the network.
4 The JMA attack
@@
As mentioned in the introduction, to avoid vanishing gradient problems, we look for a method that can be applied at the logits level. As we said, while modifying one or two logits at a time allows to carry out a close-to-optimal adversarial attack in the one-hot encoding case [7], in the more general case of DNNs based on output encoding, this approach is highly suboptimum. In this scenario, in addition to considering all the logits simultaneously, as done in the ML-C&W attack, the correlation among the logits and the effort required in the input space to move the input sample in a given direction (in the logit or feature space) must be considered. The JMA attack tackles the above problems and finds the optimal adversarial image, by solving the optimization problem in Eq. (3), under a first-order approximation of the behavior of the network function.
@@
In the following, we look at the output of the neural network function $f$  [^3] For ease of notation, in the following, we consider directly the model output $f(x)$, however the same the arguments cam be applied at the logit level. as a generic point in $\mathbb{R}^{n}$. According to Eq. (3), the optimum target point for the attacker corresponds to the point in the target decision region that can be reached by introducing the minimal perturbation in the image $x$. Such a point is not necessarily the point in the target decision region closest to the current output point $f(x)$ in the Euclidean norm, since evaluating the distance in the output space does not take into account the effort necessary to move the input sample into the desired output point. Let $x_{0}$ be the to-be-attacked image and let us denote with $\delta$ the perturbation applied by the attacker. Let $r=f(x_{0})+d$. The goal of the attacker is to determine the displacement $d$, in such a way that $r$ lies inside the decision region of the target class, and for which the Euclidean norm of the perturbation $\delta(d)$ necessary to reach $r$ is minimum, that is:
$\min_{d:\phi(f(x_{0})+d)=t}\left[\min_{\begin{subarray}{c}\delta:f(x_{0}+\delta)-f(x_{0})=d\\ x_{0}+\delta\in[0,1]^{m}\end{subarray}}\hskip 5.69046pt||\delta(d)||_{2}\right].$ (11)
@@
Without loss of generality, we will assume that all the label codewords have the same norm, and that $\phi()$ applies a minimum distance decoding rule. In this case, forcing $\phi(f(x_{0})+d)=t$ is equivalent to impose the following conditions:
@@
$\min_{\begin{subarray}{c}d:d^{T}(c_{i}-c_{t})\leq g_{ti}\\ \forall i,i=1,\cdots,l\end{subarray}}\bigg{[}\min_{\delta:f(x_{0}+\delta)-f(x_{0})=d}\hskip 5.69046pt||\delta(d)||_{2}\bigg{]}.$ (16)
Note that we neglected the constraint $x_{0}+\delta\in[0,1]^{m}$, trusting that if $\delta$ is small the constraint is always satisfied. We will then reconsider the effect of this constraint at the end of our derivation in Section 4.1.
@@
We now show that, under the first order approximation in Eq. (17), we can solve the minimization in Eq. (16), by applying the following two steps: i) we first determine the optimum perturbation moving the input sample to a given target point $r$; ii) then, we find the optimum target point $r^{*}$ by solving a constrained quadratic programming problem. Addressing i) corresponds to solving the following problem:
$\delta^{*}(d)=\arg\min_{\delta:f(x+\delta)-f(x_{0})=d}||\delta||_{2},$ (19)
@@
Given that $n<m$, under the assumption that $J$ has full rank, the solution of the minimization in Eq. (20) is:
@@
Given the above, the optimal target point $r^{*}$, and hence the optimal distance term $d^{*}$, is determined by solving the following minimization:
@@
Figure 1: Illustration of the intuition behind the formulation in (25). While $r^{\prime}$ is the point in the target region closest to $f(x_{0})$ in the Euclidean norm, $r^{*}$ is the point that can be reached with the minimum distortion in the input space.
@@
where $A=[(c_{1}-c_{t})^{T}\dots(c_{l}-c_{t})^{T}]$ has size $(l-1)\times n$, and $b$ is an $l$-long vector defined as $b_{i}=g_{ti}$. We observe that when $l>n+1$ the system is overdetermined.
@@
where $\lambda$ is a column vector with $k-1$ entries. By rewriting the objective function as $-\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+d^{T}({{(JJ^{T})}^{-1}}d+A^{T}\lambda)-b^{T}\lambda$ and after some algebra, the optimization problem can be rephrased as:
@@
The problem in (28) is a NNLS problem [20, 19], for which several numerical solvers exist [21], many of them belonging to the class of active set methods, see for instance [22, 23]. In particular, an easy-to-implement algorithm, whose complexity grows linearly with the number of label vectors, named Sequential Coordinate-Wise algorithm, has been proposed in [24].
Note that in the multi-label case, the number of label vectors grows exponentially with $n$ (since $l=2^{n}$). However, in this case, the problem can be significantly simplified, as discussed in Section 4.2.
4.1 Iterative formulation of JMA
In principle, JMA should produce an adversarial example in one-shot. In practice, however, this is not always the case, due to the fact that the linear assumption of $f$ holds only in a small neighborhood of the input $x$, hence the assumption may not be met when the distortion necessary to attack the image becomes larger. In order to mitigate this problem, we consider the following iterative version of JMA:

### v2

@@ IV Adversarial attacks against DNNs
The vulnerability of deep neural networks (DNNs) to adversarial samples has been first pointed out by Szegedy et al. in [1]. When the goal of the attacker is to induce a generic misclassification, the attack is referred to as untargeted Conversely, if the misclassification aims at a specific class, the attack is referred to as targeted. According to [1], the generation of a targeted adversarial example can be formally described as
[…]
@@
where $\delta||_{2}$ denotes the $l_{2}$ norm of the perturbation $\delta$, and $t$ is the target class of the attack. The same formulation holds for the untargeted case, with the first constraint replaced by $\phi(x+\delta)\neq\phi(x)$. The constraint $x+\delta\in[0,1]^{m}$ makes sure that the resulting adversarial example is a valid input.
IV-A Basic adversarial attacks
@@
Solving the minimization in (3) is generally hard, then the adversarial examples are determined by solving the simplified problem where a functional $\mathcal{L}({x+\delta},t)+\lambda\cdot\left\|\delta\right\|_{2}^{2}$ is minimized, where $\mathcal{L}({x+\delta},t)$ is the loss function (usually, the categorical cross entropy loss) under the target class $t$, and $\lambda$ is a parameter balancing the two terms. This problem can be solved by the box-constraint L-BFGS method [1]. Specifically, a line search is carried out to find the value of $\lambda>0$ for which the solution satisfies the adversarial condition, that is $\phi(x+\delta)=t$. It is immediate to see that, in the one-hot case, when the categorical cross entropy loss is considered, only the target class contributes to the loss term, and then the attack only cares about increasing $f_{t}(x)$, regardless of the other scores, all the more that, due to the presence of the softmax, this also implies decreasing the other outputs. To reduce the complexity of the L-BFGS attack, several suboptimal solutions have been proposed, considering the problem of minimizing the loss function $\mathcal{L}$ subject to a constraint on the perturbation $\delta$, like I-FGSM (or BIM) [12] and PGD [13].
@@
A flexible and efficient method working at the logits level has been proposed by Carlini and Wagner [6]. In the C&W attack in fact, the loss term is replaced by a properly designed function depending on the difference between the logit of the target class and the largest among the logits of the other classes. The perturbation is kept in the valid range by properly modifying the objective function. In this way, the following unconstrained problem is considered and solved (for the case of $L_{2}$ metric):
[…]
@@
where $\delta(w)=1/2(\tanh(w)+1)-x$ is the distortion introduced in the image, and the constant parameter $\kappa>0$ is used to encourage the attack to generate a high confidence attacked image. Since logits are more sensitive to modifications of the input than the probability values obtained after the softmax activation C&W attack is less sensitive to vanishing gradient problems. However, by considering only two logits at time, and neglecting the effect of the perturbation on the other logits, this approach is particularly suited to the one-hot case and is highly suboptimum when other output encoding schemes are used.
IV-B Attacks against DNNs with output encoding
@@
The adversarial attack algorithm and methods described above focus on models based on one-hot encoding. This is a favorable scenario for the attacker, who only needs to focus on increasing the output score (or logit) of the target class $t$. This is not the case with DNNs based on different output encoding schemes. The simplest examples are multi-label classifiers. From the attacker’s side, the difficulties of carrying out the attack in this case are due to the fact that the labels are not mutually exclusive, and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way. In addition, the attacker may want to modify more than one output label in a desired way. For instance, he may want to induce the classifier to interpret a daylight image showing a car driving in the rain, as a night image of a car driving with no rain. This prevents the application of the basic adversarial algorithms described in the previous section.
@@
As mentioned in Section II the problem of crafting adversarial examples against multi-label classifiers has been addressed in [9], where the C&W and the DeepFool attacks have been extended to work in a multi-label setting. The extension of C&W attack, hereafter referred to as ML-C&W, works as follows: for every output node, a hinge loss term similar to the one in Eq. (IV-A) is considered in the minimization. More specifically, the following term is added to the minimization: $\sum_{i=0}^{n}\max(0,\gamma-c_{ti}\cdot z_{i}(x+\delta))$, where $c_{t}$ is the target label vector and $\gamma$ is a confidence parameter. Note that the extension implicitly assumes that decoding is carried out element-wise. This is not true when all labels’ combinations are possible and with schemes adopting output channel coding. However, by working at the logits level, the C&W attack is general and can also be extended to work with different output encoding schemes (as discussed below).
@@
The extension of the DeepFool attack (ML-DF) works as follows: instead of targeting an objective function minimizing the distortion, ML-DF looks for the minimum perturbation that allows to enter the target region under a linear assumption on the model behavior. ML-DF works at the output score level, and hence is more prone to the vanishing gradient problem than ML-C&W. Given the threshold vector ${\rho}=({\rho}_{1},{\rho}_{2},..,{\rho}_{l})^{T}$ for the score level output ($\rho_{i}=0.5$ when $f_{i}\in[0,1]$, 0 when $f_{i}\in[-1,1]$), a vector with the distance to the boundary is derived and used to compute the perturbation. The perturbation obtained in this way is applied to the image. ML-DF applies the above procedure iteratively until the attack succeeds or the maximum number of iterations is reached. ML-DF is a greedy method and often leads to very large distortions; Moreover, the performance obtained are always inferior with respect to ML-C&W [9, 10].
@@
MLA-LP, which also implements an adversarial attack for the multi-label case, works as follows. By assuming that the loss changes are linear for small distortions, a simplified formulation of the attack is solved to minimize the $L_{\infty}$ distortion introduced in the image subject to a constraint on the loss value, requiring that the final loss is lower than the threshold value. Thanks to the adoption of the $L_{\infty}$ distortion, the problem can be easily solved by linear programming methods. Similarly to ML-DF, MLA-LP works at the final score output level. By considering the loss function instead of the model output function, MLA-LP tends to modify less the confidences of non-attacked labels with respect to ML-DF, with a lower distortion of the attacked image.
@@
[…]
@@
IV-B1 The LOTS attack
A completely different kind of attack, which can be applied to any network regardless of the output encoding scheme is the Layerwise Origin-Target Synthesis (LOTS) attack [21]. LOTS generates the adversarial example by modifying the input sample so that its representation in the feature space is as similar as possible to that of a given target sample. Formally, given the internal representation $f^{-j}(x)$ of an image $x$ at layer LjL-j and a target internal representation $f^{-j}(x_{t})$, LOTS attempts to minimize the term:
$f^{-j}(x^{t})-f^{-j}(x)||^{2}.$ (6)
@@ IV-B1 The LOTS attack
[…]
@@
IV-B2 JMA and prior art
@@
In contrast to LOTS [21], JMA defines the target point in the feature space that minimizes perturbation necessary to move the feature representation of the input sample to the target point. It does so, by relying on the Mahalanobis distance induced by the Jacobian matrix of the neural network function so to take into account the different effort, in terms of input perturbation, required to move the sample in different directions. ML-DF and MLA-LP [9, 10] are also based on the Jacobian matrix, however, in [9], the Jacobian matrix is used to implement a greedy algorithm, while in [10] is used to solve a suboptimum formulation of the attack. With JMA, instead, the original formulation of the attack by Szegedy et al. is solved under a first-order approximation. With such an approximation, in fact, the solution of the attack problem can be reduced to the solution of a constrained quadratic programming problem (where the objective function is a Mahalanobis distance term induced by the Jacobian). Moreover, our algorithm is a general one and can work either at the output level or at the logits level, regardless of the output encoding scheme used.
V The JMA attack
@@
As mentioned in Section I, our method is designed to operate at the logits level. In fact, while modifying one or two logits at a time allows to carry out a close-to-optimal adversarial attack in the one-hot encoding case [6], in the more general case of DNNs based on output encoding, this approach is highly suboptimum. In this scenario, in addition to considering all the logits simultaneously, as done in the ML-C&W attack, the correlation among the logits and the effort required in the input space to move the input sample in a given direction (in the logit or feature space) must be considered. JMA tackles the above problems and finds the optimal adversarial image, by solving the optimization problem in Eq. (3), under a first-order approximation of the behavior of the network function.
@@
In the following, we look at the output of the neural network function $f$  [^2] For ease of notation, in the following, the model output is considered, however the same procedure can be applied at the logit level. as a generic point in $\mathbb{R}^{n}$. According to Eq. (3), the optimum target point for the attacker corresponds to the point in the target decision region that can be reached by introducing the minimal perturbation in the image $x$. Such a point is not necessarily the point in the target decision region closest to the current output point $f(x)$ in the Euclidean norm, since evaluating the distance in the output space does not take into account the effort necessary to move the input sample into the desired output point. Let $x_{0}$ be the to-be-attacked image and let us denote with $\delta$ the perturbation applied by the attacker. Let $r=f(x_{0})+d$ denote a generic point in the output space. The goal of the attacker is to determine the displacement $d$, in such a way that $r$ lies inside the decision region of the target class, and for which the Euclidean norm of the perturbation $\delta(d)$ necessary to reach $r$ is minimum, that is:
[…]
The proposed attack works in two steps:
  • 1. the optimum perturbation $\delta^{*}(d)$ moving the input sample to the (generic) target point $r$ is determined;
  • 2. we find the optimum target point $r^{*}$, i.e., the optimum displacement $d^{*}$ (where $r^{*}=f(x_{0})+d^{*}$) by solving a constrained quadratic programming problem. The resulting $\delta^{*}(d^{*})$ gives the minimum perturbation necessary to reach the target region.
@@
In the following, we first rewrite the minimization in a more convenient way, then we solve Step 1 and 2 under a first order approximation. Without loss of generality, we will assume that all the label codewords have the same norm, and that $\phi()$ applies a minimum distance decoding rule. In this case, forcing $\phi(f(x_{0})+d)=t$ is equivalent to impose the following conditions:
@@
[…]
Note that we neglected the constraint $x_{0}+\delta\in[0,1]^{m}$, trusting that if $\delta$ is small the constraint is always satisfied. We will then reconsider the effect of this constraint at the end of our derivation in Section V-A.
@@
where $J_{x_{0}}$ denotes the Jacobian matrix of $f$ in $x_{0}$333ff can describe the output of the network or the logits (strictly speaking, in the latter case, $f$ should be replaced by $f^{-1}$):
We now show that the minimization in Eq. (11) can be solved under the first order approximation in Eq. (12). In particular, Step 1 corresponds to solving the following problem:
@@
Since $n<m$, under the assumption that $J$ has full rank [^5] We verified this assumption experimentally as discussed in Section VI-D., the solution of the minimization in Eq. (15) is:
@@
The optimal target point $r^{*}$, and hence the optimal distance term $d^{*}$, is then determined (Step 2) by solving the following minimization:
@@
Figure 1: Illustration of the intuition behind the formulation in (20). The point in the target region closest to $f(x_{0})$ in the Euclidean norm is $r^{\prime}$. Then, $r^{\prime}=f(x_{0})+d^{\prime}$, where $d^{\prime}$ is the minimum norm displacement necessary to reach the target region. Instead, the point that can be reached with the minimum distortion in the input space, i.e., the point which minimizes $\delta$, is $r^{*}$ (this point corresponds to a displacement $d^{*}$ having larger norm than $d^{\prime}$). $r^{*}$ is the optimal target point that the attack wants to reach.
@@
where $A=[(c_{1}-c_{t})^{T}\dots(c_{l}-c_{t})^{T}]$ has size $(l-1)\times n$, and $b$ is an (l1)(l-1)-long vector defined as $b_{i}=g_{ti}$. We observe that when $l>n+1$ the system is overdetermined.
@@
where $\lambda$ is a column vector with l1l-1 entries. By rewriting the objective function as $\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+d^{T}({{(JJ^{T})}^{-1}}d+A^{T}\lambda)-b^{T}\lambda$ and after some algebra, the optimization problem can be rephrased as:
@@
The problem in (23) is a NNLS problem [26, 25], for which several numerical solvers exist [27], many of them belonging to the class of active set methods, see for instance [28, 29]. In particular, an easy-to-implement algorithm, whose complexity grows linearly with the number of label vectors, named Sequential Coordinate-Wise algorithm, has been proposed in [30]. Theorem 1 identifies the minimum perturbation necessary to enter the target region $\delta^{*}(d^{*})$ (simply indicated as $\delta^{*}$ in the following). Then the adversarial image is computed as $x_{adv}=x+\delta^{*}$.
Note that in the multi-label case, the number of label vectors grows exponentially with $n$ (since $l=2^{n}$). However, in this case, the problem can be significantly simplified, as discussed in Section V-B.
V-A Iterative formulation of JMA
In principle, JMA should produce an adversarial example in one-shot. In practice, however, this is not always the case, due to the fact that the linear assumption of $f$ holds only in a small neighborhood of the input $x$, hence the assumption may not be met when the distortion necessary to attack the image is larger. In order to mitigate this problem, we consider the following iterative version of JMA: